An object weighing 296 N in air is immersed
in water after being tied to a string connected
to a balance. The scale now reads 265 N .
Immersed in oil, the object appears to weigh
272 N .
Find the density of the object.
Answer in units of kg/m3.
Don't round answer.

Find the density of the oil.
Answer in units of kg/m3.
1
Expert's answer
2020-02-05T12:27:57-0500

"W-W_w=\rho_wgV"

"296-265=1000(9.8)V"

"V=0.00316\ m^3"

Thus,


"\rho_{object}=\frac{W}{gV}"

"\rho_{object}=\frac{296}{(0.00316)(9.8)}=9560\frac{kg}{m^3}"

"W-W_o=\rho_ogV"

"296-272=\rho_o(9.8)(0.00316)"

"\rho_{o}=775\frac{kg}{m^3}"